a population size N = 150 has μ = 8 and standard deviation of 5.4. What is the probability that a random size n=20 will have a mean of 9.5 above


1
Expert's answer
2021-05-07T14:18:40-0400

Given,

Mean of population "\mu=8"

Standard deviation "\sigma = 5.4"

Using

"P(\bar X >9.5)= \dfrac{\bar X-\mu}{\frac{\sigma}{\sqrt n}}"


"P( \bar X>9.5)= \dfrac{9.5-8}{\frac{5.4}{\sqrt{20}}}=\dfrac{1.5}{1.20}=1.25"


So, "P(Z>1.25)=0.1056"





Hence, Probability P(Z>1.25)= 0.1056